OpenStudy (anonymous):

* I will fan & medal! xoxo. * :D The staff at the community center make a rectangular ice skating rink every year. They use 140 meters of boarding to enclose it. So the area of the rink depends upon the width undefined chooses. This rule describes how the area depends on the width, w: w(70 ??? w) or 70w ??? w^2 Width | Area 10 | 600 15 | 825 20 | 1000 25 | ? The table shows some of the values for this function. Help complete the able. If the width is 25 metes, what will the area be? <3

OpenStudy (anonymous):

*table

OpenStudy (solomonzelman):

I am sturggling to understand the rule they are going by...

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

Well, if you know that the width is 25 meters, then use: A=l•w You know that P=2(l+w) So you can use that to find the length. 140=2(l+25) solve for length

OpenStudy (solomonzelman):

So find l, and then the Area

OpenStudy (anonymous):

l = 45

OpenStudy (solomonzelman):

Yes, correct

OpenStudy (solomonzelman):

And then: A=w•l A=25•45 A=?

OpenStudy (anonymous):

1,125?

OpenStudy (solomonzelman):

I don't have a calc. so lets do that.... 25•45= 25•(25+20)= 25•25+25•20= 625+500= 1125. Yes:)

OpenStudy (anonymous):

Yay! thank you :) <3

OpenStudy (solomonzelman):

I like practicing every technique that I can encounter while doing a question, this is why I have done what I did just now....

OpenStudy (solomonzelman):

By the way, in Russia for example as long as other Ueropean countries, when you use a comma you indicate a decimal so I try to use space (source: from Engineering).

OpenStudy (solomonzelman):

e.g \(1{\tiny~~}125\)

OpenStudy (solomonzelman):

in any case, i know you meant a thousand, tho'

OpenStudy (anonymous):

Oh I see! Thank you for informing me! ^_^

OpenStudy (solomonzelman):

Don't worry about it tho now:) You are welcome, and good luck!

OpenStudy (triciaal):

agree with the above just wanted to add the function A = (70 - w)w

OpenStudy (solomonzelman):

Yes, that is true:)

OpenStudy (solomonzelman):

By the way good choice of meters, as they are the "SI Units"... don't want to overwhelm on the same question tho'..... cu